Building the New Dome

Daily events and images of the BBSO dome replacement project.

29 November Early today the first shipment of new dome parts arrived from MFG/Ratech. The parts include the new dome base ring, part of the shutter and some lower dome panels. Some tests of the dome anchor bolts must be done before assembly can begin.
Meanwhile we are moving the new earthshine dome out to its foundation.
01 December Today the new dome base ring is being put into place. The ring is made of four sections which are bolted together. The base ring also has the bottom parts of the new arch girders welded onto it.
05 December Today the first course of dome panels was finished.
06 December The MFG/Ratech crew is working on shore today. They are preparing the dome arch girders and shutter track for installation. The preparation work is expected to take several days.
13 December The second row of dome panels is being installed today. These are the panels with the vent louvers.
14 December More second-row panels being installed. Here you can see the motors to control the vents and the anti-bird grills installed on the inside of the louvers.
18 December The arch girders are now both in place, and the third course of dome panels is being installed. Also the fixed panels between the arch girders have been installed.
20 December The last of the third row of panels are now up. The MFG/Ratech crew is now attaching a tarp to cover the shutter opening and the top of the dome. The crew will return next month to finish the job.
21 December The MFG/Ratech crew is securing a tarp on the dome to cover the top and shutter opening. The crew will return next month to install the dome top panels, finish the caulking and install the dome shutter, aperture cover and motors.
23 December More pictures showing the tarp the MFG/Ratech crew placed on the dome.
31 December It seems all of our plans have come to naught. The tarp was blown off by high winds on Monday the 26th. There is a storm expected tonight and another one Monday. Jeff and Randy have arranged tarps inside the dome to direct the water into the basement, from where it can flow into the lake. All of the computers in the dome have been turned off in case water does reach them.
03 January We had a storm yesterday, with rain, snow and high winds. The tarp system mostly worked well, but there were a few leaks. The MFG/Ratech crew has returned to resume work on the dome. They are starting by drilling the bolt holes in the top panels.
04 January The MFG/Ratech crew is installing one of the top panels on the dome.
12 January The MFG/Ratech crew is installing the first shutter panel. The telescope will look through this aperture at the Sun.
13 January The remaining shutter panels were installed today. Next job is to install the motors for the shutter and the dome rotation. Today we tested one of the shutter drive motors in the lab.
02 February The exterior of the dome is complete except for the door over the aperture and the skirt to cover the dome base ring. Conduit is being installed inside the dome to carry power and control signals to the louvers The motors for the shutter and dome rotation and the azimuth drive chain have not yet been installed. Also changes to the drive chains for the shutter need to be made.

© BBSO/NJIT - Last modified: March 2, 2006.